OpenStudy (anonymous):

solve for the following"v": e= m v^2/2

OpenStudy (anonymous):

use algebra

OpenStudy (anonymous):

2e=mv^2

OpenStudy (anonymous):

how do i get v alone?

OpenStudy (anonymous):

you have to continue with the algebra sequence

OpenStudy (anonymous):

2e/m=v^2

OpenStudy (anonymous):

theres one more step

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

youre welcome

OpenStudy (goformit100):

First of all @hello_sunshine : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct